OpenStudy (anonymous):

integral x sqrt (3x^2 + 4) dx

OpenStudy (anonymous):

try u = 3x^2

OpenStudy (anonymous):

alright thanks

OpenStudy (anonymous):

i mean u = 3x^2 + 4, my bad

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

can you explain how to use u substitution for this problem

OpenStudy (anonymous):

\[\int\limits_{}^{}x \sqrt{3x^{2} + 4} dx\] let u = 3x^2 + 4 du = 6x dx dx = 1/(6x) du so... \[\frac{ 1 }{ 6 } \int\limits_{}^{} \sqrt{u} du\]

OpenStudy (anonymous):

did I do this correctly?\[\frac{ (3x^2 + 4)^{3/2} }{9}+c\]

OpenStudy (anonymous):

never mind

OpenStudy (anonymous):

yeah you did it right